What is Henry’s age when told that he is 5 times as old as his son, but will in 4 years time be only 3 times as old?
Answer by stanbon(75887)

What is Henry’s age when told that he is 5 times as old as his son, but will in 4 years time be only 3 times as old?
---------------------
Equations:
h = 5s
h+4 = 3(s+4)
---
Substitute for "h" and solve for "s":
5s+4 = 3s+12
2s = 8
s = 4 (Son's age now)
---
Solve for "h":
h = 5s
h = 5*4 = 20 (Henry's age now)
=============================
Cheers,
Stan H.
===========